OpenStudy (bbb911):

Find all numbers for which the rational expression is not defined 8/2x

OpenStudy (anonymous):

real numbers i think

OpenStudy (anonymous):

@wio

jimthompson5910 (jim_thompson5910):

if the expression is \[\Large \frac{8}{2x}\] then you cannot divide by zero This means the denominator 2x cannot be equal to zero. If it were, then 2x = 0 2x/2 = 0/2 x = 0 So if x = 0, then the denominator is zero. That means x = 0 makes the expression undefined.

OpenStudy (bbb911):

oh i see thanks jim ! :3

jimthompson5910 (jim_thompson5910):

you're welcome